Circular No. 9132

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

CBAT Director: Daniel W. E. Green; Room 209; Dept. of Earth and Planetary Sciences; Harvard University; 20 Oxford St.; Cambridge, MA 02138, U.S.A.

cbat@iau.org  cbatiau@eps.harvard.edu  ISSN 0081-0304

                URL http://www.cfa.harvard.edu/iau/cbat.html

COMET C/2010 F4 (MACHHOLZ) Don Machholz reports his visual discovery of a diffuse comet with

coma diameter ∼ 2′ (DC = 2), using a 0.47-m f/4.8 reflector (77×) at his home in Colfax, CA, U.S.A., in the course of his comet-hunting program on Mar. 23, as tabulated below; no motion was detected in 20 minutes due to encroaching twilight. Machholz found the comet again of similar appearance (the magnitude then given as 11.5–12) on the morning of Mar. 26, indicating that it is moving rapidly towards the sun. Machholz had searched for 607 hours since his previous find on 2004 Aug. 27 (cf. IAUCs 8394, 8395). After posting on the Minor Planet Center’s ‘NEOCP’ webpage, three Japanese amateurs have reported confirmation of the new comet. Y. Ikari (Moriyama, Shiga-ken, 0.26-m reflector + CCD) reports a coma diameter of 0′.5. K. Kadota (Ageo, Saitama-ken, 0.25-m reflector + CCD) finds that his images taken through clouds show the object to be diffuse (coma diameter 1′.5) with central condensation and no tail. K. Yoshimoto (Yamaguchi-ken, 0.25-m reflector) reports that a visual observation yields total mag 11.0 and coma diameter 1′.3 (DC = 3) on Mar. 26.830 UT.

2010 UT  α_{2000}  δ_{2000}  Mag. Observer
Mar. 23.526  22^{h}54.^{m}2  +31^{o}22′  11  Machholz
26.515  23 23.5  +32 14  12  _{”}
26.80750 23 25 47.20 +32 13 14.8 13.9 Ikari
26.81116 23 25 49.37 +32 13 16.8  _{”}
26.81482 23 25 51.44 +32 13 19.0  _{”}
26.82056 23 25 54.80 +32 13 22.6  _{”}
26.82158 23 25 55.40 +32 13 23.3 12.3 Kadota
26.82296 23 25 56.18 +32 13 23.6  _{”}
26.82734 23 25 58.73 +32 13 26.3 14.7 Ikari

The following parabolic orbital elements are extremely uncertain: T = 2010 Apr. 10.669 TT, q = 0.58698 AU, ω = 135^{o}.062, Ω = 245^{o}.662, i = 86^{o}.831 (equinox 2000.0). Ephemeris using H_{7.5} = 12.0:

2010 TT α_{2000} δ_{2000} Δ r ϵ β Mag.
Mar. 25 23^{h}08.^{m}35 +31^{o}50.^{′}0 1.237 0.693 34.^{o}1 53.^{o}7 11.3
26 23 17.99 +32 04.4 1.241 0.682 33.3 53.4 11.2
27 23 27.63 +32 15.1 1.246 0.671 32.5 53.0 11.2
28 23 37.26 +32 22.1 1.251 0.661 31.7 52.5 11.1
29 23 46.84 +32 25.3 1.258 0.651 30.9 51.9 11.1
30 23 56.35 +32 24.8 1.266 0.642 30.1 51.3 11.1
31 0 05.76 +32 20.6 1.274 0.633 29.3 50.5 11.0
2010 March 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green